OpenStudy (anonymous):

\[\int\limits_{}^{}(\cos x- \cos 2x) dx/1-cosx\]

OpenStudy (anonymous):

\[=\int\limits_{}^{}((\cos x-1)+(1-\cos 2x))dx/(1-\cos x)\] \[=\int\limits_{}^{}dx+\int\limits_{}^{}2\sin ^{2}xdx/(1-\cos x)\] \[x+\int\limits_{}^{}2\sin ^{2}xdx/(2\sin ^{2}x/2)\] \[x+\int\limits_{}^{}4\cos ^{2}x/2dx\]i think now you can do the simple integration

OpenStudy (anonymous):

it should be -x and how did you get this?? \[\int\limits\limits_{}^{}4\cos ^{2}x/2dx\]

OpenStudy (anonymous):

yes it should be -x. sin x=2sin(x/2)cos(x/2). use this in the gration.